OpenStudy (anonymous):

(x-1/x^2)/(x-1/x^3) = solve please!

OpenStudy (ybarrap):

We can reduce this not solve because there is nothing on the right-hand side. Simplify the following: (x-1/x^2)/(x-1/x^3) Put each term in x-1/x^2 over the common denominator x^2: x-1/x^2 = x^3/x^2-1/x^2: (x^3/x^2-1/x^2)/(x-1/x^3) x^3/x^2-1/x^2 = (x^3-1)/x^2: ((x^3-1)/x^2)/(x-1/x^3) x^3-1 = x^3-1^3: ((x^3-1^3)/x^2)/(x-1/x^3) Factor the difference of two cubes. x^3-1^3 = (x-1) (x^2+x 1+1^2): ((x-1) (x^2+x+1^2)/x^2)/(x-1/x^3) 1^2 = 1: (((x-1) (x^2+x+1))/x^2)/(x-1/x^3) Put each term in x-1/x^3 over the common denominator x^3: x-1/x^3 = x^4/x^3-1/x^3: ((x-1) (x^2+x+1))/(x^2 x^4/x^3-1/x^3) x^4/x^3-1/x^3 = (x^4-1)/x^3: ((x-1) (x^2+x+1))/(x^2 (x^4-1)/x^3) x^4-1 = (x^2)^2-1^2: ((x-1) (x^2+x+1))/(x^2 (x^2)^2-1^2/x^3) Factor the difference of two squares. (x^2)^2-1^2 = (x^2-1) (x^2+1): ((x-1) (x^2+x+1))/(x^2 (x^2-1) (x^2+1)/x^3) x^2-1 = x^2-1^2: ((x-1) (x^2+x+1))/(x^2 ((x^2-1^2) (x^2+1))/x^3) Factor the difference of two squares. x^2-1^2 = (x-1) (x+1): ((x-1) (x^2+x+1))/(x^2 ((x-1) (x+1) (x^2+1))/x^3) Multiply the numerator by the reciprocal of the denominator, ((x-1) (x^2+x+1))/(x^2 ((x-1) (x+1) (x^2+1))/x^3) = ((x-1) (x^2+x+1))/x^2×x^3/((x-1) (x+1) (x^2+1)): ((x-1) (x^2+x+1) x^3)/(x^2 (x-1) (x+1) (x^2+1)) ((x-1) (x^2+x+1) x^3)/(x^2 (x-1) (x+1) (x^2+1)) = (-1+x)/(-1+x)×((x^2+x+1) x^3)/(x^2 (x+1) (x^2+1)) = ((x^2+x+1) x^3)/(x^2 (x+1) (x^2+1)): ((x^2+x+1) x^3)/(x^2 (x+1) (x^2+1)) Combine powers. ((x^2+x+1) x^3)/(x^2 (x+1) (x^2+1)) = ((x^2+x+1) x^(3-2))/((x+1) (x^2+1)): ((x^2+x+1) x^3-2)/((x+1) (x^2+1)) 3-2 = 1: Answer: ((x^2+x+1) x)/((x+1) (x^2+1))